what is the functional value requested in the problem:

F(x)=x^2-t^2

a.) F(t)=
b.) F(5t)=

Answer by jim_thompson5910(35256)

a.) F(t)=

F(x)=x^2-t^2

F(t)=t^2-t^2

F(t)=0t^2

F(t)=0
-------------------------------------------------------
b.) F(5t)=


F(x)=x^2-t^2

F(5t)=(5t)^2-t^2

F(5t)=25t^2-t^2

F(5t)=24t^2


--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------